### v1

# Sub-wavelength optical lattice in 2D materials

Abstract
Recently, light-matter interaction has been vastly expanded as a control tool for inducing and enhancing many emergent non-equilibrium phenomena [1, 2]. However, conventional schemes for exploring such light-induced phenomena [3, 4, 4, 5, 6, 7, 8, 9, 10, 11, 12] rely on uniform and diffraction-limited free-space optics, which limits the spatial resolution and the efficiency of light-matter interaction. Here, we overcome these challenges using metasurface plasmon polaritons (MPPs) to form a sub-wavelength optical lattice. Specifically, we report a “nonlocal” pump-probe scheme where MPPs are excited to induce a spatially modulated AC Stark shift for excitons in a monolayer of MoSe2, several microns away from the illumination spot. Remarkably, we identify nearly two orders of magnitude reduction for the required modulation power compared to the free-space optical illumination counterpart. Moreover, we demonstrate a broadening of the excitons’ linewidth as a robust signature of MPP-induced periodic sub-diffraction modulation. Our results open new avenues for exploring power-efficient light-induced lattice phenomena below the diffraction limit in active chip-compatible MPP architectures.
.1 Main
Recent decades have witnessed a surge of interest in using light as a powerful tool to induce, control, and enhance exotic emergent non-equilibrium phenomena in quantum and topological materials [1, 2]. For example, light has been widely utilized for inducing AC Stark shift [3, 4, 5], Floquet bands [6, 7], signatures of magnetization [8, 9], superconductivity [10], anomalous Hall effect [11], and melting gapped state into semi-metallic ones [12]. Common schemes for exploring such light-induced phenomena are based on diffraction-limited free-space optical spectroscopic techniques, which do not provide spatial patterns to imprint sub-wavelength lattice features [13]. These methods fundamentally put a bound on the spatial resolution and the efficiency of light-matter interaction due to the limited confinement of free-space optical beams. Moreover, pump-probe spectroscopy of emergent light-induced phenomena typically is performed “locally”, i.e., the pump and the probe illuminate and investigate the same spot. Such local pump-probe techniques cause unavoidable parasitic thermal and pump-induced noise, which necessitates complex heat management and pump-rejection methods [14]. The development of novel platforms and spectroscopy techniques to address these challenges remains an active area of research [15].
Concurrently, the emergence of emitter-integrated nano-plasmonics has enabled strong sub-wavelength confinement of optical fields for efficient light-matter interaction. Of particular interest are localized and propagating surface plasmon polaritons (SPPs), which are excitations at the metal-dielectric interface whose electromagnetic field decays exponentially with distance from the surface, offering remarkable control at near-field [16, 17, 18]. For example, active nano-plasmonic devices have been recently used for strong light-matter coupling [19, 20, 21, 22, 23], probing dark excitons and cascaded energy transfer [24, 25] and directional emission [26]. On the downside, these devices offer limited propagation length and functionalities, which motivated the development of metasurface plasmon polaritons (MPPs) [27, 28, 29] with superior performance. MPPs have enabled several interesting effects such as Purcell-enabled optical modulators [30], exciton manipulation [31], enhanced harmonic generation [32], and plasmon-exciton strong coupling [33, 34], which are explored with local optical spectroscopy.
Here, we design and realize a novel MPP-based nonlocal pump-probe scheme for inducing sub-wavelength optical lattices. Specifically, the excited MPPs imprint a one-dimensional lattice for excitons, resulting from a spatially modulated AC Stark shift.
Figure 1: Sub-diffraction nonlocal pump-probe spectroscopy of excitons. (a) Scheme for inducing AC Stark shift on a MoSe2 monolayer using metasurface plasmon polaritons (MPPs) in a 1D lattice of silver grooves. The periodically modulated MPP modes are excited with a pump laser (red beam) at the top coupler (TC). The MPP device is designed to strongly confine the plasmonic modes, which propagate in the y-direction of the lattice with negligible diffraction within $700-800$ nm, which is close to the excitonic resonance of $\sim 750$ nm in MoSe2. A secondary broadband laser (white beam) probes the MoSe2 excitons. (b) Schematic of the device cross-section, including the hBN-encapsulated MoSe2 on the silver grooves. The periodicity of the lattice is $a=150$ nm, the height is $h=160$ nm, and the width is $w=90$ nm. The red curve shows the strong $|E|^{2}$ modulation at the position of the MoSe2 monolayer for a pump of 800 nm. Excitons present at positions of peak electric field intensity observe larger blue shifts compared to the ones near the center of the groove. (c) Normalized electric field intensity $|E|^{2}$ at the x-z cross-section of the device. The position of the TMD is marked by the dashed horizontal line, which is 10 nm above the surface. (d) Iso-frequency contours of the MPP device for different wavelengths of MPPs showing suppressed diffraction in the range 700-800 nm. (e-g) Finite-difference time-domain (FDTD) simulations comparing SPP propagation loss over a distance of 3 $\mu$m on a 2D sheet of silver, a single groove waveguide, and the proposed MPP device at 800 nm respectively. In (e-g) a dipole source placed 10 nm above the surface is used to excite the SPPs.
.2 Concept
Figure 1(a) illustrates our scheme. The metasurface device consists of a 1D array of silver nano-grooves, and is designed to support a 1D lattice of MPPs with negligible diffraction around wavelengths close to the excitonic resonance. A pump laser (marked in red) is coupled into the MPP modes through the top coupler. The traveling MPPs form a sub-diffraction periodic potential for the MoSe2 monolayer, whose exciton resonance lies at a wavelength $\sim 750$ nm. A broadband beam (marked in white) probes the modulation of the transition metal dichalcogenide (TMD) excitons at a spot distant from the top coupler.
Figure 1(c) shows the normalized electric field intensity $|E|^{2}$ in the x-z plane (shown only for one groove) at 800 nm obtained from finite-difference time-domain (FDTD) simulations. The electric field intensity is maximum at hot spots near the sharp corners of the air-silver interface and decays exponentially with distance from the surface; this is a characteristic feature of SPPs. The TMD is located $\sim 10$ nm above the surface, as marked by the dashed white line in Fig. 1(b). The red curve shows the electric field intensity $|E|^{2}$ at the monolayer’s position. Importantly, due to this sub-diffraction variation of $|E|^{2}$, the excitons experience a position-dependent AC stark shift: those located at the local maxima of $|E|^{2}$ experience a larger blue shift compared to other excitons (represented by the dark and light blue ellipses in Fig. 1(b); see SI for detailed study of spatial dependence of $|E|^{2}$ and the full scheme of the experimental setup).
The diffractionless propagation of MPPs is a fundamental advantage for our nonlocal AC Stark shift scheme. To understand this feature, we calculate the dispersion of the device using the FDTD method. Figure 1(d) shows the iso-frequency contours of the device for different wavelengths. Here, $k_{x}$ and $k_{y}$ are the wavevectors along the x and y directions. Notably, at wavelengths of $\sim 700-800$ nm, the MPPs propagate with maximally suppressed diffraction. Moreover, for lower wavelengths (blue lines) the dispersion is hyperbolic with a negative angle of refraction, whereas it is elliptical with a positive angle of refraction for longer wavelengths (red line). Our numerical simulations closely follow the analytical dispersion relationship of the MPP device given by the coupled waveguide formalism under a slowly-varying amplitude approximation [35, 36]: $k_{y}=k_{wg}+2t\cos(k_{x}a)$, where $k_{wg}$ is the propagation constant of the MPP waveguide lattice, and $t$ is the tunneling amplitude between nearest-neighbor grooves in the lattice (see SI for details).
Moreover, the choice of an MPP device structure over other conventional plasmonic platforms such as 2D sheets and single grooves provides a noticeable advantage in long-distance power delivery due to their diffractionless propagation. To confirm this, the propagation of the excitation for each case is shown in Fig. 1(e-g). Over a propagation length of 3 $\mu$m, the MPP device has $\sim 5$ and $\sim 50$ folds enhanced power transmission compared to its single groove, and 2D sheet counterparts, respectively. Details of the simulations and a more comprehensive analysis of the wavelength-dependent operation of our device can be found in the SI.
.3 Device fabrication and characterization
Figure 2: MPP device fabrication and characterization. (a) High-resolution scanning electron microscope (SEM) images of representative fabricated MPP device. A 1D array of grooves is terminated with a top coupler (TC) and a bottom coupler (BC) to in-couple and out-couple the propagating MPPs. The top left inset shows a close-up view of the TC and the grooves. The bottom left inset shows a perspective view with a cross-section of the grooves. (b) The measured transmission spectrum (normalized by the broadband input laser spectrum) of the MPP device, shows a transmission peak around the pump wavelength of 800 nm. The inset shows the simulated transmission spectrum. (c) The spatial pattern of the emission from the output coupler for three different wavelengths: 770 nm, 796 nm and 828 nm. The red dashed line shows the position of the pump laser. (d) A false-color microscope image of the device after stacking the MoSe2 monolayer. TC, BC and the MPP device are marked by dashed boxes.
Figure 2(a) shows a high-resolution scanning electron microscopy (SEM) image of a representative MPP device, which is fabricated on a 500 nm thick silver layer sputtered on a silicon carbide substrate. Each device consists of 40 silver grooves (each 20 $\mu$m long) and has a top coupler, and a bottom coupler on each end of the metasurface. The top and bottom left insets show the top coupler-MPP device interface, and a cross-section of the waveguides, respectively. To prevent the silver from oxidation, the chip is coated with a 5 nm thick layer of Al2O3. Further details of the fabrication are available in the SI.
We begin by characterizing the transmission spectrum of the device before stacking the TMD on the top, by coupling a broadband laser into one of the couplers and measuring the output from the other. The results are shown in Fig. 2(b). The transmission is maximum for a wavelength of $\sim 800$ nm and it decreases as a function of the detuning, in agreement with the simulated iso-frequency contours (Fig. 1(d)). The inset shows the simulated normalized transmission spectrum for the device monitored after 3 $\mu$m propagation of the MPPs. At greater detuning from the peak at $\sim 800$ nm, the measured transmission is smaller than in the simulation, since at large diffraction angles the MPPs do not successfully reach the bottom coupler. The asymmetric nature of the normalized transmission can be attributed to the longer propagation lengths for the elliptical MPPs [28]. To experimentally characterize the diffraction in the MPP device, we measure the spatial pattern of the emission from the output coupler for three different wavelengths as shown in Fig. 2(c). The red line shows the position of the pump laser at the input coupler. We observe diffracted outputs at wavelengths of 770 nm and 828 nm, whereas there is minimal diffraction at 796 nm, which is very close to our designed pump wavelength. Simulation and details of this wavelength-dependent behavior are available in the SI. After the initial MPP device characterizations, we complete the integrated device by stacking the hBN-encapsulated MoSe2 monolayer on the MPP device. Figure 2(d) shows a false-color microscope image after the stacking procedure.
.4 Nonlocal and local AC Stark shift measurements
Figure 3: Nonlocal modification of exciton energy with MPPs. (a) Schematic of our nonlocal AC Stark shift measurement scheme that utilizes a sub-diffraction spatially-modulated MPP pump. The device is pumped at the top coupler to generate MPP modes in the metasurface. The MoSe2 is probed $\sim 3\mu$m away from the pump. (b) Nonlocal AC Stark shift was measured in device D1 while the delay between the pump and probe pulses was varied. The white lines show the extracted exciton peak wavelength from numerical fitting. (c) The extracted change in linewidth $\Delta\Gamma$ of the exciton extracted from the numerical fitting of the nonlocal AC Stark shift data. (d) Schematic of traditional local AC Stark shift measurement scheme that utilizes a diffraction-limited Gaussian optical pump. (e) Local AC Stark shift measured in device D2 (MoSe2 on silver substrate). (f) The extracted change in linewidth $\Delta\Gamma$ of the exciton extracted from the numerical fitting of the local AC Stark shift data. The error bars represent one standard deviation error in parameter estimation from numerical fitting.
To demonstrate the sub-diffraction nonlocal AC Stark shift scheme introduced in Fig. 1, we excite the MPP modes using a 150 fs pulsed laser at 800 nm wavelength, as shown schematically in Fig. 3(a). The pump excites the MPPs in several grooves that lie under our diffraction-limited pump (see the SI for a discussion on the multi-groove excitation). They travel along the y-direction of the device and modulate the MoSe2 monolayer located on top of the MPP device. Importantly, the excited MPPs have a spatial periodic intensity profile with sharp features below the diffraction limit (see SI for simulations of the mode profiles). To probe the AC stark shift induced by the MPPs, we perform absorption spectroscopy using a weak pulsed broadband laser placed 3 $\mu$m away from the pump in a reflection configuration. We note that the pump is red-detuned by $\sim 50$ nm from the excitonic energy to prevent photoluminescence.
Figure 3(b) shows the normalized reflectivity spectrum as a function of pump-probe temporal delay $\Delta t$. Positive $\Delta$t refers to the probe reaching the sample before the pump, and vice-versa. We observe an AC Stark shift of $\sim$0.65 nm at $\Delta t=0$. The white line shows the extracted exciton wavelength from the numerical fitting (see the SI for the detail of the numerical fitting). In the data presented in Fig. 3(b), we pump the top coupler with 100 mW of power to create the MPPs. Taking into account the estimated efficiency of the device, this translates to an “MPP pump” power of 0.33 mW at the position of the TMD under the probe. The complete analysis of the power efficiency, as well as pump power and pump detuning dependence of the AC Stark shift, is presented in the SI.
.4.1 Power efficiency comparison
Next, we confirm the superior power efficiency of our nonlocal pump-probe scheme, which arises from MPP-induced strong field confinement, by considering the optical power required to generate the same AC Stark shift using the conventional local pump-probe counterpart. In the local case, the pump and the probe are spatially overlapped as shown in Fig. 3(d). We perform the local AC Stark shift measurement on a TMD placed on an unpatterned silver substrate (device D2) as shown in Fig. 3(e). We observe an identical AC shift of $\sim 0.65$ nm using $\sim 25$ mW of pump power. Remarkably, when compared to the MPP-induced nonlocal AC Stark shift with an effective pump power of 0.33 mW, the nonlocal AC Stark shift scheme requires $\sim 90$ times less pump power to observe a comparable AC Stark shift for the same detuning (see the SI for details). Numerically, we estimate that the maximum field enhancement at the position of the TMD for our nonlocal scheme can be as high as 130 times the diffraction-limited optical pump intensity (see SI for detailed calculation). The slight discrepancy between the experimental observation and numerical estimate is likely due to fabrication imperfections like surface roughness, rounded groove corners, and deviation of the TMD position from its nominal value in the z-direction. Moreover, this remarkable observed enhancement is further substantiated by measuring a similar nonlocal AC Stark shift effect on a different device, which highlights the robustness of our scheme (see the data for device D4 in the SI). Furthermore, we confirm that the MPP-enabled strong field confinement is central to our scheme compared to 2D sheets (discussed earlier in Fig. 1(e-g)). To achieve this, we perform the same nonlocal scheme on a similar TMD device placed on a 2D sheet. Using an identical pump power strength, no AC Stark shift is observed in this case (see the nonlocal measurement of device D2 in the SI).
.4.2 Sub-diffraction modulation of excitons
We demonstrate another intriguing optical lattice-induced effect, in the form of an excitonic linewidth broadening, presented in Fig. 3(c). The excitons under our diffraction-limited probe laser experience a strong spatial variation of the electric field intensity of the MPPs as shown in Fig. 3(a). This periodic sub-diffraction modulation of electric field intensity (which can vary as strongly as $\sim 50$% as shown in the inset of Fig. 3(a)) manifests as a linewidth broadening of $\sim 1$ nm at $\Delta t=0$. This observation is in sharp contrast with the local counterpart, where there is no noticeable change in the linewidth at $\Delta t=0$ (Fig. 3(f)) (for a similar observation and discussion for a TMD monolayer placed on an unpatterned silica substrate, see the data for the device D3 in the SI).
We further confirm that this linewidth broadening originates from the sub-diffraction spatial modulation of the MPP by performing local AC Stark shift measurements on the top of the MPP device D1 (see Fig. S17). We observe a comparable linewidth broadening to the nonlocal case arising from the sub-diffraction modulation provided by the device, albeit, using a much larger pump power than the nonlocal case. Moreover, we note that thermal effects, which take place in much slower timescales than our short 150 fs pulse duration, are not likely to play a significant role in this observation. For the pump powers and detunings used in our experiments, we can also rule out possible contributions from power broadening (see the SI for detailed analysis).
.5 Summary and outlook
In summary, we demonstrated a novel TMD-integrated plasmonic metasurface approach for efficient modulation of excitons below the diffraction limit. Due to its intrinsic nonlocal pump-probe character [37], our platform provides a useful protocol for spatial on-chip pump rejection as well as suppressing parasitic thermal effects [14]. More generally, our results suggest that lattice SPPs in MPP device architectures can become a new and versatile toolbox to actively manipulate matter excitations below the diffraction limit.
Beyond our exemplary AC Stark shift case, a wide range of other light-induced phenomena can be investigated below the diffraction limit with superior power efficiency. This includes light-induced magnetization [8], superconductivity [38], quantum Hall effects [11], electronic polarization [39], polariton modulators [40] and plasmonic effects in van der Waals heterostructures [41, 42]. Another intriguing direction is the sub-wavelength periodic modulation of materials and matter excitations harnessing the versatile tunability and controllability of our MPPs. An example includes optical imprinting of lattices [13], where the optical drive forms a lattice structure in space. The lattice nature of MPPs also may enable engineering chiral light-matter interaction [31] using plasmonic spin Hall states [28].
In addition to light-induced phenomena, our platform could be useful for cavity-induced physics [2, 43, 44, 18]. In particular, the strong spatial confinement of light in such plasmonic structures combined with the hyperbolic photonic modes of the lattice can lead to intriguing phenomena such as meditated long-range interactions between distant emitters [45, 46].
.6 Methods
Device fabrication: The single-crystalline silver films (500 nm thick) and buffer layer platinum film were deposited by DC sputtering on a SiC wafer, similar to the process reported in Ref. [28]. Next, the devices were patterned on the silver by electron beam lithography with an Al2O3 hard mask, followed by reactive ion etching and plasma etching. After etching, another thin layer of Al2O3 is deposited on the devices to prevent the silver from oxidizing. After initial characterization of the plasmonic devices, the hBN-MoSe2-hBN heterostructure was transferred onto the device using the standard dry transfer method with PDMS (Polydimethylsiloxane) and PC (Polycarbonate) stamps. Complete details of the device fabrication are available in the SI.
Experimental setup: The sample was mounted in a Montana cryostat at 4 K. For reflectivity measurements, we used a confocal microscopy setup. A Ti:Sapphire laser tuned at 800 nm, with a pulse duration of 150 fs and a repetition rate of 80 MHz was used as the pump. The pump beam was split with a polarizing beam splitter, and one branch of it was sent into a supercontinuum generation non-linear fiber to create a broadband probe pulse. Meanwhile, the other branch was sent to a motorized delay stage. This allowed us to control the relative path length difference between the pump and the probe pulses up to a precision of 100 nm. Next, the pump and the generated broadband probe beams were sent through an optical chopper operating at 1 kHz. This was to prevent heating the sample while having a high pump peak intensity. The chopper was placed at the focal spot of a telescope such that both pump and probe were focused on the same spot at the chopper. A half-waveplate linear polarizer combination was used to cross-polarize the residual pump before detection of the signal in a spectrometer with a 300 grooves per mm diffraction grating, and a CCD camera. A detailed schematic of the experimental setup is available in the SI.
FDTD simulations: The electromagnetic field profiles and iso-frequency contour simulations were performed using the FDTD method in Lumerical. A dipole placed 10 nm above the surface was used to excite the MPP modes of interest in the simulations, unless otherwise specified. The propagating MPPs were monitored using frequency domain monitors to extract quantities like propagation loss, coupling efficiency to gratings, field confinement, far-field radiation pattern, etc. The iso-frequency contours were calculated by performing Fourier transformations of the simulated field profiles in the x-y plane over a large area.
.7 Acknowledgements
The authors wish to acknowledge fruitful discussions with Aaron Sternbach, Zubin Jacob, Andrey Grankin, Deric Session, and Nick Martin. This research was supported by DARPA HR00112490310, the National Science Foundation DMR-2145712, and the Department of Energy DE-SC-0022885 grant. This research used Quantum Material Press (QPress) of the Center for Functional Nanomaterials (CFN), which is a US Department of Energy, Office of Science User Facility, at Brookhaven National Laboratory under contract no. DE-SC0012704. K.W. and T.T. acknowledge support from the JSPS KAKENHI (grant nos. 20H00354, 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan, for hBN synthesis.
.8 Competing interests
The authors declare no competing interests.
.9 Data availability
All of the data that support the findings of this study are reported in the main text and Supplementary Information. Source data are available from the corresponding authors on reasonable request.